OpenStudy (anonymous):

Someone please help me! Problem: A 2.00 kg ball (ball 1), moving to the right at a velocity of +4.00 m/s on a frictionless table, collides head-on with a stationary 7.70 kg ball (ball 2). Find the final velocities of both balls if the collision is a completely inelastic. (I already found the values for the final velocities if the collision is elastic; ball 1=-2.35 and ball 2=1.65)

OpenStudy (anonymous):

in inelastic collision: \[M _{1}V_{1}+M _{2}V _{2}=V(M _{1}+M _{2})\] \[2.0kg \times4.0m/s+0.0m/s \times 7.7kg = V(2.0kg+7.7kg)\] \[8kg m/s \div 9.7kg= 0.825m/s\]